Two cars are approaching one another. Car A moves
Answer #1

4) Applying Doppler shift formula we get

f'=\frac{c+v_o}{c-v_s}\times f \quad \quad \text{c=speed of sound,} v_o=\text{speed of observer},v_s=\text{speed of source} \\ \implies f'=\frac{c+\frac{c}{4}}{c-\frac{c}{6}}\times1000\\ \implies f'=1500Hz

5)Flux is \Phi=\vec{B}\cdot\vec{A}=Ft

Induced EMF is \varepsilon=-\frac{d\Phi}{dt}=-F

6)Magnetic field is

\frac{\mu_0}{2}\left( \frac{I}{a}-\frac{I'}{b}\right )\\ =\frac{\mu_0}{2}\left( \frac{20}{0.1}-\frac{5}{0.2}\right )\\ =1.099\times10^{-4}T